OpenStudy (anonymous):

Simplify the square root of 48. 6 the square root of 3 4 the square root of 3 3 the square root of 6 3 the square root of 4

OpenStudy (anonymous):

Is it A?

OpenStudy (anonymous):

not likely first off it is never A secondly see if you can factor 48 as a perfect square times some other number

OpenStudy (anonymous):

you might see that \(48=3\times 16\) making \[\sqrt{48}=\sqrt{16\times 3}=\sqrt{16}\sqrt3=4\sqrt3\]

OpenStudy (gk99):

yes B

OpenStudy (anonymous):

so it's b